find the value of sin1*sin2*sin3*...*sin179
(note: all angles are in degrees)

Grade:

3 Answers

Aman Bansal
12 years ago

Dear Student,

The general product is

     P[n] = Product[Sin[k Pi/n], {k,1,n-1}] = n/2^(n-1),

and the special case is 

     P[180] = 180*2^(-179) 

To prove this, let

     v = Exp[Pi i/n]

be a primitve (2n)th root of unity.  Then we have

     Sin[k Pi/n] = (v^k - v^-k)/(2i) 

                 = (1 - v^(-2k))(v^k/(2i)).

Therefore

     P[n] = Product[1 - v^(-2k), {k,1,n-1}] Product[v^k/(2i), {k,1,n-1}].

But the first product is simply the factorization of

     (z^(2n) - 1)/(z^2 - 1) = z^(2n-2) + z^(2n-4) + ... + 1

with z = 1.  Consequently, this first product has value n.  The
second product is found by summing the exponents, giving

     P[n] = n v^Sum[k, {k,1,n-1}]/(2i)^(n-1)

          = n v^((n-1)n/2) /(2i)^(n-1)

          = n Exp[Pi i (n-1)/2]/(2 Exp[Pi i/2])^(n-1)

          = n/2^(n-1) Exp[Pi i (n-1)/2 - Pi i(n-1)/2]

          = n/2^(n-1),

Best Of luck

Plz Approve the answer...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….
So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Abhijat Pandey
12 years ago

The answer is ZERO!!

As the series is in increasing order... therefore, there is sin90° in between... which is 0 and multiplying with 0 is alway 0...

plz approve my answer!! :)

Kaushik Abhay Mudgal
6 years ago
The answer will be 0 as it will include sin 90, which is zero and anything multiplied to zero is zero.